OpenStudy (anonymous):

Solve: y=3x+5 x-y=-3

OpenStudy (anonymous):

your second equation can be rewritten as x=y-3. substitute that into your first equation y=3x+5 y=3(y-3)+5 y=3y-9+5 y=3y-4 -2y=-4 y=2 then substitute 2 into your equation for y and solve for x 2=3x+5 -3=3x x=-1

OpenStudy (anonymous):

Thank You rorytsm!!!

OpenStudy (anonymous):

no problem